09 March 2020 UT SOUL-C Night 1 Notes

Closed at sunset for clouds and wind

Wanting to check pupil and FS alignment once again, I took a pupil image, but it looks blurry and the 3 screws are not visible. First images were with the ARGOS swing arm in, but even when it was out, it was blurry.

Aha... once we took the N375 field image, we saw right away that the guide probe was on-axis. The field image taken with the probe on-axis allowed at least to see that the N30_FieldStop was pretty well centered at 1033, 1031 (0494)

04:00 Still Closed for clouds.

Closed dome tests of ESM mode.

Alfio making some SW changes to remove some steps in setting up, closing the loop in full-AO mode that are unecessary for ESM mode. This wasn't working and we moved on.

Check/measure pupil wobble.

Measure position of lens that compensates for wobble (due to misalignment between instrument and re-rorator, K-mirror) as the LUCI rotator is moved from 0 to 360. Rather than a nice ellipse, its motion traces a 'jerky' sort of ellipse. But it is within the limits.

At bin 4 (vs bin1), looks smoother. But it is close to a lower limit. Fourth one at apparently the same settings yields a different sort of shape - not too far from elliptical, but not as nice as the previous one. Questions about this.

NCPA measurements.

Plan is to do modes 5 and 6 first, then 5 through 10.

Set a 256x256 subwin to contain the star at all PAs (considering pupil wobble)
Set readmode = cube

Doug running a script to collect NCPA data.
  • At each rotator angle, take a series of LUCI images first with z5 coefficient stepping from low to high, centered at zero, and then stepping through z6. Find the coeff which minimizes the Strehl.
  • 07:37 AO arbitrator froze - data collection stalled for about 30 min with rotator at 180. Resumed. FYI, Doug's data-taking script ~lucifer/lcsp/bin/commitReadout.sh
  • 08:26 loop opened - from 180 to 210 deg. A lot went to zero when restarting the arbitrator, including re-rotator.
  • 08:41 z stage was not set. Restart at 210 deg.
Arcetri going to run tests while Doug analyzes data.

On-Sky

11:17 Opening. RH finally started to drop and roof looks clear.

ARGOS cal unit parked.

12:00 Focus/Collimate on little wfs. Seeing is >2".

12:07 Going to a R=8.9 star to close the loop.

12:23 Launched M3 script. PA=46 deg, PAR=72, ROT=341

12:26. About 8 stars in the 15" diameter porthole on CCD47.
  • re-rotator error. Didn't go to 341 but instead to -18.
  • Manually moved LUCI1 rotator so it is above zero and resent the preset.
  • first image taken, 0806 (H), but before gains were optimized (start AO clicked before opt gain). But then run AO failed.
    • FWHM ~ 35 pix (imexam "r") or 0.55". (FWHM from Aladin is 15.5 pix or 0.23" ???)
  • Doug and Shane stopping AO
  • Resent script. Starting with 0807. PSF gradient over the field. Images on right elongated diagonally, 814 stars are doubled. 816 not doubled. Enrico identifies it as an isoplanatic effect, though not all the stars really seem to be pointing to the AO ref - there seems to be a slight deviation from this. Seeing was about 1.5-2" during this time.
13:07 Too bright. Loops opened.

-- OlgaKuhn - 09 Mar 2020
Topic revision: r6 - 09 Mar 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback